Kyky232:

Find the total surface area of a cylinder with a height of 5 cm and radius of 2 cm. Leave your answer in terms of π. 28π cm2 24π cm2 70π cm2 45π cm2

Kyky232:

its supposed to be cm^2

snowflake0531:

Surface area formula \[2(\pi r^2) + 2\pi r(h)\] plug the values in \[2(4\pi) + 4\pi(5)\] Can you simplify that

Kyky232:

uh i got 87 sumn

Kyky232:

wait

snowflake0531:

Keep it in pi form

Kyky232:

I got 45 this time

snowflake0531:

can you tell me how you got there

Kyky232:

I put 2(4π)+4π(5) into a calculator

snowflake0531:

don't use a dumb calculator use your brain --- also, keep it in pi form though

Kyky232:

My brain isnt smart and idk how to do this

Kyky232:

a calculators smarter than me

snowflake0531:

you don't know 2 times 4? you dont know 4 times 5? sheesh okay

snowflake0531:

\(2(4\pi) + 4(5\pi) = 8\pi + 20\pi = 28 \pi^2\)

snowflake0531:

*pi cm^2

Kyky232:

so its A?

snowflake0531:

positive